15 postcards, 10 envelopes, and a notepad cost 1 dollar and 68 cents. The envelope is 8 times cheaper than the notepad and 2 cents more expensive than the postcard. What is the price of the postcard, the envelope, and the notepad?
Answer by ikleyn(52803)

.
15 postcards, 10 envelopes, and a notepad cost 1 dollar and 68 cents. The envelope is 8 times cheaper than the notepad
and 2 cents more expensive than the postcard. What is the price of the postcard, the envelope, and the notepad?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let x = postcard.
Then envelope = x+2
     notepad  = 8*(x+2)

x + (x+2) + 8*(x+2) = 168.

10x + 18 = 168  --->  10x = 150  --->  x = 15.

Answer.  Postcard = 15 cents, envelope = 17 cents, notepad = 8*17 cents = $1.36.

Lesson to learn from this solution: this problem is for ONE unknown.

Not for three.